The length of the base of an isosceles triangle is 2 centimemters shorter than the length of either of the other sides. If the perimeter of the triangle is 16 centimeters, find the length of each side of the triangle.
Answer by checkley75(3666)

b+2s=16
b=s-2 thus
s-2+2s=16
3s=16+2
3s=18
s=18/3
s=6 answer for the length of the sides.
proof
(6-2)+2*6=16
4+12=16
16=16